bag of smokeless fuel are labeled as weighing 100kg but actual weights are normally distributed with a mean of 102kg and a standard deviation of 4kg.
What proportion of bag will weigh more than 106kg
1
Expert's answer
2020-05-19T20:17:48-0400

Let "X=" the weight of bag: "X\sim N(\mu, \sigma^2)"


"Z={X-\mu\over \sigma}\sim N(0,1)"

Given "\mu=102, \sigma=4"


"P(X>106)=1-P(X\leq106)=1-P(Z\leq{106-102\over 4})="

"=1-P(Z\leq1)\approx0.1587"

"15.87\ \%"